OpenStudy (anonymous):

"Easy" Calculus! Two similar questions: 1) If f(4) = 5 and f'(4) = 2/3, then calculate the derivative of f inverse of 5. 2) If g(7) = 3 and g'(3) = 5/6 and g'(7) = 3/4. then calculate the derivative of f inverse of 3 The answer to #1 is 3/2 and #2 is 4/3, but could someone fill me in on the intermediate steps/general method for solving?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\left(f^{-1}(x)\right)'=\frac{1}{f'(f^{-1}(x)}\]

OpenStudy (anonymous):

and so \[\left(f^{-1}(5)\right)'=\frac{1}{f'(f^{-1}(5)}\]

OpenStudy (anonymous):

now for your problem \[f(4)=5\implies f^{-1}(5)=4\]

OpenStudy (anonymous):

so \[\left(f^{-1}(5)\right)'=\frac{1}{f'(f^{-1}(5)}\]\[=\frac{1}{f'(4)}=\frac{1}{\frac{2}{3}}=\frac{3}{2}\]

OpenStudy (anonymous):

second one is similar

OpenStudy (anonymous):

both these problems were to get you to practice \[\left(f^{-1}(x)\right)'=\frac{1}{f'(f^{-1}(x)}\]

OpenStudy (anonymous):

Okay, so I've see the \[(f '^ {-1}(x))\] equals 1/f' whatever..., but was wondering, when I'm solving for the derivative of f inverse of 5, its ok just to use 1 over f'(4)?

OpenStudy (anonymous):

Oh wait. I see why.